# Spiking Neural Networks for Continuous Control via End-to-End Model-Based Learning

### v2

@@ Abstract
Despite recent progress in training spiking neural networks (SNNs) for classification, their application to continuous motor control remains limited. Here, we demonstrate that fully spiking architectures can be trained end-to-end to control robotic arms with multiple degrees of freedom in continuous environments. Our predictive-control framework combines Leaky Integrate-and-Fire dynamics with surrogate gradients, jointly optimizing a forward model for dynamics prediction and a policy network for goal-directed action. We evaluate this approach on both a planar 2D reaching task and a simulated 6-DOF Franka Emika Panda robot. Results show that SNNs can achieve stable training and accurate torque control, establishing their viability for high-dimensional motor tasks. An extensive ablation study highlights the role of initialization, learnable time constants, and regularization in shaping training dynamics. We conclude that while stable and effective control can be achieved, recurrent spiking networks remain highly sensitive to hyperparameter settings, underscoring the importance of principled design choices.
@@ 1 Introduction
Despite these challenges, spiking networks have already proven highly effective in sensory processing, classification, and unsupervised learning (Neftci et al.,, 2019; Yin et al.,, 2021; Zenke and Ganguli,, 2018; Rossbroich et al.,, 2022; Tavanaei et al.,, 2019), where their sparse, event-driven nature and internal state dynamics provide concrete advantages. These models, inspired by how biological neurons communicate, are increasingly adopted in abstract computational settings for their ability to process temporally structured input and operate with remarkable efficiency. Yet most practical machine learning, especially in continuous control, remains dominated by deep learning architectures with continuous-valued activations.
@@
Research on spiking control spans diverse methodologies, learning approaches, and tasks, with multiple surveys synthesizing progress to date. While these studies have advanced understanding of adaptive neuron models, network architectures, bio-plausibility, local learning rules, and various hardware constraints, they often focus on aspects somewhat peripheral to scalable, general-purpose control. Several overviews emphasize the growing relevance of robotics applications (Bing et al.,, 2018; Lanillos et al.,, 2021; DeWolf,, 2021; Oikonomou et al.,, 2025), while methodological surveys cover deep-learning-inspired SNN learning and training advances (Tavanaei et al.,, 2019; Neftci et al.,, 2019; Zenke and Vogels,, 2021; Eshraghian et al.,, 2023). Together, these establish SNN control as promising but technically challenging, especially in high-dimensional continuous domains.
@@
One of the earliest approaches embeds analytical control equations directly into spiking networks. Slijkhuis et al., (2023) developed closed-form spike coding controllers, Agliati et al., (2025) introduced predictive spiking control for linear systems, and Traub et al., 2021a extended this paradigm with recurrent spiking networks for dynamic action inference. Such controllers offer elegance and interpretability, drawing inspiration from control theory, but typically rely on one-step least-squares optimization rather than learning. As a result, they depend heavily on accurate system models and often lack robustness to unmodeled dynamics in real-world environments.
@@
Moving beyond purely analytical formulations, a large body of work has investigated bio-inspired local plasticity. Fernández et al., (2021) implemented STDP-like adaptation for robotic arm reaching, Juarez-Lora et al., (2022) applied R-STDP to arm control with changing friction, and Chen et al., (2020) developed a bio-inspired controller for a 4-DoF arm. Building on the Neural Engineering Framework (NEF), DeWolf et al., (2016), Iacob et al., (2020), and Marrero et al., (2024) applied NEF-based models to arm control, linking adaptive computation, cognition, and robotics, with later Loihi-based extensions to 7-DoF arms (DeWolf et al.,, 2023). Beyond arm reaching, Schmidgall and Hays, (2023) proposed synaptic motor adaptation for locomotion, while Jiang et al., (2025) introduced a fully spiking locomotion controller. These approaches make use of data-driven adaptation and offer some biological plausibility, but they remain limited to relatively low-complexity tasks and do not yet scale to high-DoF continuous control.
@@
Another important direction focuses on ANN-to-SNN conversion and neuromorphic deployment. Zanatta et al., (2023) demonstrated spiking DRL agents optimized for neuromorphic accelerators, highlighting energy efficiency benefits. Zhao et al., (2020) implemented closed-loop control on the iCub robot using DYNAP-SE, Paredes-Vallés et al., (2024) realized fully neuromorphic drone flight, and DeWolf et al., (2023) deployed Loihi-based control of a 7-DoF arm. These studies show the feasibility of deploying spiking control policies on neuromorphic hardware, but they are typically constrained by hardware-specific neuron models and architectural limitations, often prioritizing efficiency over generality.
@@
In contrast to hardware-driven or conversion-based methods, surrogate gradient techniques (Neftci et al.,, 2019; Zenke and Vogels,, 2021) enable true end-to-end optimization in spiking domains by approximating spike derivatives, bringing deep-learning-style training into SNNs. Originally developed for classification, these methods have since been adapted to control, with stabilization strategies such as fluctuation-driven initialization further improving training dynamics (Rossbroich et al.,, 2022). Applied to reinforcement learning, Tang et al., (2021) used population coding for MuJoCo tasks, Chen et al., (2024) proposed noisy spiking actors to enhance exploration, Park et al., (2025) extended TD3 to 3D arm control, and Oikonomou et al., (2023) applied deep RL to 6-DoF reaching. Beyond arms, Kumar et al., (2025) developed a spiking DQN for mobile robot navigation, Traub et al., 2021b demonstrated recurrent SNNs scaling to many-joint arms, and Zanatta et al., (2024) explored PPO-based DRL integration in robotic tasks. Together, these works highlight scalability and flexibility, overcoming many limitations of earlier local or conversion approaches. However, most efforts focus on policy optimization alone, without predictive modeling components.
@@
A particularly promising next step is the integration of learned predictive models into the control loop, not only for state estimation, but as generative world models capable of producing synthetic experience for policy learning (Ha and Schmidhuber,, 2018; Hafner et al.,, 2025; Huebotter et al.,, 2022). This paradigm, already transformative in the ANN-based reinforcement learning community, could help SNN controllers scale to high-dimensional, long-horizon tasks without prohibitively expensive data collection. In the SNN domain, several first steps exist: Taniguchi et al., (2023) reviewed predictive coding for cognitive robotics, Agliati et al., (2025) provided a formal account of predictive spiking control, and Zhu et al., (2024) applied SNNs to autonomous driving by combining perception and planning. Most closely aligned with model-based RL, Capone and Paolucci, (2024) proposed dreaming in recurrent SNNs, though their approach remains limited to simpler settings. Together, these works suggest that predictive modeling may help scale SNN controllers, but the joint, end-to-end optimization of predictive and control networks in fully spiking architectures for high-DoF continuous arm control remains underexplored—a gap directly addressed by the present work.
@@
Grouping by task domains further illustrates this methodological progression. For arm reaching, analytical formulations such as closed-form spike coding and predictive control have been explored (Slijkhuis et al.,, 2023; Agliati et al.,, 2025), followed by local plasticity and NEF-inspired approaches (Chen et al.,, 2020; Juarez-Lora et al.,, 2022; Fernández et al.,, 2021; DeWolf et al.,, 2016; Iacob et al.,, 2020; Marrero et al.,, 2024). Neuromorphic deployments have extended these ideas to real-world settings, with Loihi-based 7-DoF arms (DeWolf et al.,, 2023) and iCub controllers implemented on DYNAP-SE (Zhao et al.,, 2020). More recently, surrogate-gradient-driven RL has scaled SNN controllers to high-DoF arms, including 3D reaching with TD3 (Park et al.,, 2025), 6-DoF hybrid RL control (Oikonomou et al.,, 2023), and recurrent architectures spanning many-joint systems (Traub et al., 2021b, ). Beyond arms, locomotion has been approached with adaptive rules (Schmidgall and Hays,, 2023; Jiang et al.,, 2025), drones with neuromorphic controllers (Paredes-Vallés et al.,, 2024), and autonomous driving with predictive frameworks (Zhu et al.,, 2024). Broader contributions include surveys synthesizing the field (Bing et al.,, 2018; Lanillos et al.,, 2021; Oikonomou et al.,, 2025), methodological advances in training (Tavanaei et al.,, 2019; Eshraghian et al.,, 2023), and stabilization techniques for scalable optimization (Rossbroich et al.,, 2022). Finally, predictive-model-based approaches remain in their infancy, with only initial steps toward integrating generative world models into spiking control (Capone and Paolucci,, 2024; Taniguchi et al.,, 2023).
@@ 2.1 Pred-Control SNN
Our SNN architecture is depicted in Figure 1. It draws inspiration from world models Ha and Schmidhuber, (2018) and predictive coding approaches to robot control Lanillos et al., (2021); Taniguchi et al., (2023) and is composed of: $i)$ a prediction network $\bm{\upsilon}$, which learns to predict the robot’s future state evolution as a function of its current state and actions; and ii)ii) a policy network $\bm{\pi}$, which computes the best action (continuous control signals) to apply depending on the state of the system and a given end-effector target position. Breaking up the network into two parts allows us to use the prediction model $\bm{\upsilon}$ to simulate state trajectories during learning of the policy model $\bm{\pi}$. We can directly propagate the gradients of the distance between the predicted future positions and the target (policy loss) with respect to the policy network parameters through the prediction network.
All neurons inside the networks are modeled as leaky integrate-and-fire (LIF) units with optional adaptive thresholds (ALIF). These neuron models introduce several time constant parameters $\tau$, which require careful tuning, as they control the temporal dynamics of information integration and eligibility traces for surrogate-gradient-based backpropagation. Details of the LIF and ALIF implementations, including time constants and reset mechanisms, are described in detail in subsection 2.2 and Appendix I. For each environment time step, the SNNs are running for a set of internal sub-steps (e.g., 7) to propagate input information through the network with spike signals for sufficient temporal resolution. All networks and learning dynamics have been implemented using our Control Stork framework [^1] https://github.com/jhuebotter/control_stork (based on PyTorch) and the specific experiments for this project are available on GitHub [^2] https://github.com/jhuebotter/spiking_control. All relevant hyperparameters for the final model are summarized in Appendix O. A brief overview of the notation for the following sections is given in Table 1.
@@
The prediction network $\bm{\upsilon}$ comprises two spiking LIF populations and one non-spiking leaky integrator readout population. The first spiking layer includes full recurrent connectivity. Network size and architectural variants were explored empirically (see Appendix D). Its input consists of the current robot state $\bm{s}_{t}$ (joint angles, end-effector pose) and control action $\bm{u}_{t}$ (joint velocity or acceleration). The network outputs a predicted state change $\Delta\hat{\bm{s}}_{t}$, which is added to the previous estimate $\hat{\bm{s}}_{t}$ to yield $\hat{\bm{s}}_{t+1}$. The prediction error $e_{\hat{s}}$ at time $t$ is measured by
@@
and is averaged over a trajectory of length $T$. All learnable parameters are summarized as $\bm{\theta_{\upsilon}}$, yielding the compact notation $\bm{\upsilon}(\hat{\bm{s}}^{\prime}|\bm{s},\bm{u},\bm{\theta_{\upsilon}})$. For implementation and training details, refer to Appendix F.
@@
where $K^{*}$ refers to the subset of state dimensions describing the target. All learnable parameters are denoted by $\bm{\theta_{\pi}}$, and we write the policy model as $\bm{\pi}(\bm{u}|\bm{s},\bm{s}^{*},\bm{\theta_{\pi}})$. See Appendix G for details.
@@
The scalar input vector $\bm{x}_{t}$ ($\bm{s}_{t},\bm{u}_{t}]$ for prediction network, $\bm{s}_{t},\bm{s}^{*}_{t}]$ for policy network) is linearly transformed by
@@
with input weights $W_{\text{in}}$ and bias $\bm{b}_{\text{in}}$, which serves as the input current $I_{\text{inj},t}$ to the first spiking layer.
To obtain the continuous-valued output vector $\bm{y}_{t}$ (actions or predictions), we use the membrane potentials $U_{t}$ of a non-spiking layer. These first and final weight matrices are scaled differently compared to other weight parameters in the network to generate stable internal dynamics as well as output on the correct expected magnitude. The output scaling also differs for prediction and control models, as the predicted changes in state $\Delta\hat{\bm{s}}(t)$ are typically much smaller than the control signals $\bm{u}(t)$. In the policy network, a $\tanh$ activation function ensures that control outputs remain within [1,1][-1,1].

### v3

@@ Abstract
Despite recent progress in training spiking neural networks (SNNs) for classification, their application to continuous motor control remains limited. Here, we demonstrate that fully spiking architectures can be trained end-to-end to control robotic arms with multiple degrees of freedom in continuous environments. Our predictive-control framework combines Leaky Integrate-and-Fire dynamics with surrogate gradients, jointly optimizing a forward model for dynamics prediction and a policy network for goal-directed action. We evaluate this approach on both a planar 2D reaching task and a simulated 6-DOF Franka Emika Panda robot with torque control. In direct comparison to non-spiking recurrent baselines trained under the same predictive-control pipeline, the proposed SNN achieves comparable task performance while using substantially fewer parameters. An extensive ablation study highlights the role of initialization, learnable time constants, adaptive thresholds, and latent-space compression as key contributors to stable training and effective control. Together, these findings establish spiking neural networks as a viable and scalable substrate for high-dimensional continuous control, while emphasizing the importance of principled architectural and training design.
@@ 1 Introduction
Despite these challenges, spiking networks have already proven highly effective in sensory processing, classification, and unsupervised learning (Neftci et al., 2019; Yin et al., 2021; Zenke and Ganguli, 2018; Rossbroich et al., 2022; Tavanaei et al., 2019), where their sparse, event-driven nature and internal state dynamics provide concrete advantages. These models, inspired by how biological neurons communicate, are increasingly adopted in abstract computational settings for their ability to process temporally structured input and operate with remarkable efficiency. Yet most practical machine learning, especially in continuous control, remains dominated by deep learning architectures with continuous-valued activations.
@@
Research on spiking control spans diverse methodologies, learning approaches, and tasks, with multiple surveys synthesizing progress to date. While these studies have advanced understanding of adaptive neuron models, network architectures, bio-plausibility, local learning rules, and various hardware constraints, they often focus on aspects somewhat peripheral to scalable, general-purpose control. Several overviews emphasize the growing relevance of robotics applications (Bing et al., 2018; Lanillos et al., 2021; DeWolf, 2021; Oikonomou et al., 2025; Mompó Alepuz et al., 2024), while methodological surveys cover deep-learning-inspired SNN learning and training advances (Tavanaei et al., 2019; Neftci et al., 2019; Zenke and Vogels, 2021; Eshraghian et al., 2023; Oikonomou et al., 2025; Yamazaki et al., 2022). Together, these establish SNN control as promising but technically challenging, especially in high-dimensional continuous domains.
@@
One of the earliest approaches embeds analytical control equations directly into spiking networks. Slijkhuis et al. (2023) developed closed-form spike coding controllers, Agliati et al. (2025) introduced predictive spiking control for linear systems, and Traub et al. (2021a) extended this paradigm with recurrent spiking networks for dynamic action inference. Such controllers offer elegance and interpretability, drawing inspiration from control theory, but typically rely on one-step least-squares optimization rather than learning. As a result, they depend heavily on accurate system models and often lack robustness to unmodeled dynamics in real-world environments.
@@
Moving beyond purely analytical formulations, a large body of work has investigated bio-inspired local plasticity. Fernández et al. (2021) implemented STDP-like adaptation for robotic arm reaching, Juarez-Lora et al. (2022) applied R-STDP to arm control with changing friction, and Chen et al. (2020) developed a bio-inspired controller for a 4-DoF arm. Complementary to plasticity-driven controllers, compact bio-inspired neuromorphic SNN controllers have also been proposed to produce smooth reaching dynamics for multi-DoF arms (Polykretis et al., 2023b). Building on the Neural Engineering Framework (NEF), DeWolf et al. (2016), Iacob et al. (2020), and Marrero et al. (2024) applied NEF-based models to arm control, linking adaptive computation, cognition, and robotics, with later Loihi-based extensions to 7-DoF arms (DeWolf et al., 2023). Beyond arm reaching, Schmidgall and Hays (2023) proposed synaptic motor adaptation for locomotion, while Jiang et al. (2025) introduced a fully spiking locomotion controller. These approaches make use of data-driven adaptation and offer some biological plausibility, but they remain limited to relatively low-complexity tasks and do not yet scale to high-DoF continuous control.
Another important direction focuses on ANN-to-SNN conversion/transfer. Representative conversion pipelines include converting trained networks for robotic inverse kinematics and control (Volinski et al., 2022), converting RL policies for event-based obstacle avoidance (Salvatore et al., 2020), and converting discrete-action DQN policies to spiking Q-networks (Patel et al., 2019).
@@
A closely related direction is the deployment of spiking controllers on neuromorphic hardware. As an example of directly trained spiking DRL agents targeting neuromorphic accelerators, Zanatta et al. (2023) demonstrated spiking models optimized for neuromorphic accelerators, highlighting energy efficiency benefits. Zhao et al. (2020) implemented closed-loop control on the iCub robot using DYNAP-SE, Paredes-Vallés et al. (2024) realized fully neuromorphic drone flight, and DeWolf et al. (2023) deployed Loihi-based control of a 7-DoF arm. Deployment-focused studies show the feasibility of deploying spiking control policies on neuromorphic hardware, but they are typically constrained by hardware-specific neuron models and architectural limitations, often prioritizing efficiency over generality.
@@
In contrast to hardware-driven or conversion-based methods, surrogate gradient techniques (Neftci et al., 2019; Zenke and Vogels, 2021) enable true end-to-end optimization in spiking domains by approximating spike derivatives, bringing deep-learning-style training into SNNs. Originally developed for classification, these methods have since been adapted to control, with stabilization strategies such as fluctuation-driven initialization further improving training dynamics (Rossbroich et al., 2022). Applied to reinforcement learning, Tang et al. (2021) used population coding for MuJoCo tasks, Chen et al. (2024) proposed noisy spiking actors to enhance exploration, Park et al. (2025) extended TD3 to 3D arm control, and Oikonomou et al. (2023) applied deep RL to 6-DoF reaching. Beyond arms, Kumar et al. (2025) developed a spiking DQN for mobile robot navigation, Traub et al. (2021b) demonstrated recurrent SNNs scaling to many-joint arms, and Zanatta et al. (2024) explored PPO-based DRL integration in robotic tasks. Together, these works highlight scalability and flexibility, overcoming many limitations of earlier local or conversion approaches. However, most efforts focus on policy optimization alone, without predictive modeling components.
@@
A particularly promising next step is the integration of learned predictive models into the control loop, not only for state estimation, but as generative world models capable of producing synthetic experience for policy learning (Ha and Schmidhuber, 2018; Hafner et al., 2025; Huebotter et al., 2022; Stöckl et al., 2024). This paradigm, already transformative in the ANN-based reinforcement learning community, could help SNN controllers scale to high-dimensional, long-horizon tasks without prohibitively expensive data collection. In the SNN domain, several first steps exist: Taniguchi et al. (2023) reviewed predictive coding for cognitive robotics, Agliati et al. (2025) provided a formal account of predictive spiking control, and Zhu et al. (2024) applied SNNs to autonomous driving by combining perception and planning. Beyond our control setting, recent neuromorphic-motivated work has shown that local prediction learning can support planning in high-dimensional domains (Stöckl et al., 2024), that cognitive-map structure combined with neural sampling can enable goal-directed imagination and planning (Lin et al., 2025), and that similar cognitive-map learners can be applied to mapless navigation under edge constraints (Polykretis and Danielescu, 2024). Most closely aligned with model-based RL, Capone and Paolucci (2024) proposed dreaming in recurrent SNNs, though their approach remains limited to simpler settings. Together, these works suggest that predictive modeling may help scale SNN controllers, but the joint, end-to-end optimization of predictive and control networks in fully spiking architectures for high-DoF continuous arm control remains underexplored—a gap directly addressed by the present work.
@@
Grouping by task domains further illustrates this methodological progression. For arm reaching, analytical formulations such as closed-form spike coding and predictive control have been explored (Slijkhuis et al., 2023; Agliati et al., 2025), followed by local plasticity and NEF-inspired approaches (Chen et al., 2020; Juarez-Lora et al., 2022; Fernández et al., 2021; DeWolf et al., 2016; Iacob et al., 2020; Marrero et al., 2024; Volinski et al., 2022). Related spiking joint-control motifs have also been explored for articulated multi-joint motion generation outside classical robotics, e.g., via evolutionarily selected spiking controllers for realistic 3D motion synthesis (Polykretis et al., 2023a). Neuromorphic deployments have extended these ideas to real-world settings, with Loihi-based 7-DoF arms (DeWolf et al., 2023) and iCub controllers implemented on DYNAP-SE (Zhao et al., 2020). More recently, surrogate-gradient-driven RL has scaled SNN controllers to high-DoF arms, including 3D reaching with TD3 (Park et al., 2025), 6-DoF hybrid RL control (Oikonomou et al., 2023), and recurrent architectures spanning many-joint systems (Traub et al., 2021b). Beyond arms, locomotion has been approached with adaptive rules (Schmidgall and Hays, 2023; Jiang et al., 2025), drones with neuromorphic controllers (Paredes-Vallés et al., 2024; Salvatore et al., 2020), mobile robot navigation (Polykretis and Danielescu, 2024), and autonomous driving with predictive frameworks (Zhu et al., 2024). Broader contributions include surveys synthesizing the field (Bing et al., 2018; Lanillos et al., 2021; Oikonomou et al., 2025), methodological advances in training (Tavanaei et al., 2019; Eshraghian et al., 2023), and stabilization techniques for scalable optimization (Rossbroich et al., 2022). Finally, predictive-model-based approaches remain in their infancy, with only initial steps toward integrating generative world models into spiking control (Capone and Paolucci, 2024; Taniguchi et al., 2023).
@@ 2.1 Pred-Control SNN
Our SNN architecture is depicted in Figure 1. It draws inspiration from world models Ha and Schmidhuber (2018) and predictive coding approaches to robot control Lanillos et al. (2021); Taniguchi et al. (2023) and is composed of: $i)$ a prediction network $\bm{\upsilon}$, which learns to predict the robot’s future state evolution as a function of its current state and actions; and ii)ii) a policy network $\bm{\pi}$, which computes the best action (continuous control signals) to apply depending on the state of the system and a given end-effector target position. Breaking up the network into two parts allows us to use the prediction model $\bm{\upsilon}$ to simulate state trajectories during learning of the policy model $\bm{\pi}$. We can directly propagate the gradients of the distance between the predicted future positions and the target (policy loss) with respect to the policy network parameters through the prediction network.
All neurons inside the networks are modeled as leaky integrate-and-fire (LIF) units with optional adaptive thresholds (ALIF). These neuron models introduce several time constant parameters $\tau$, which require careful tuning, as they control the temporal dynamics of information integration and eligibility traces for surrogate-gradient-based backpropagation. Details of the LIF and ALIF implementations, including time constants and reset mechanisms, are described in detail in Section 2.2 and Appendix I. For each environment time step, the SNNs are running for a set of internal sub-steps (e.g., 7) to propagate input information through the network with spike signals for sufficient temporal resolution. All networks and learning dynamics have been implemented using our Control Stork framework [^1] https://github.com/jhuebotter/control_stork (based on PyTorch) and the specific experiments for this project are available on GitHub [^2] https://github.com/jhuebotter/spiking_control. All relevant hyperparameters for the final model are summarized in Appendix R. A brief overview of the notation for the following sections is given in Table 1.
@@
The prediction network $\bm{\upsilon}$ comprises two spiking LIF populations and one non-spiking leaky integrator readout population. The first spiking layer includes full recurrent connectivity. Network size and architectural variants were explored empirically (see Appendix D). Its input consists of the current robot state $\bm{s}_{t}$ (joint angles, end-effector pose) and control action $\bm{u}_{t}$ (joint velocity or acceleration). The network outputs a predicted state change $\Delta\hat{\bm{s}}_{t}$, which is added to the previous estimate $\hat{\bm{s}}_{t}$ to yield $\hat{\bm{s}}_{t+1}=\hat{\bm{s}}_{t}+\Delta\hat{\bm{s}}_{t}$. The prediction error $e_{\hat{s}}$ at time $t$ is measured by
@@
where $K$ denotes the dimensionality of the full state vector $\bm{s}_{t}$. During learning this error further is averaged over a trajectory of length $T$. All learnable parameters are summarized as $\bm{\theta_{\upsilon}}$, yielding the compact notation $\bm{\upsilon}(\hat{\bm{s}}_{t+1}|\bm{s}_{t},\bm{u}_{t},\bm{\theta_{\upsilon}})$. For implementation and training details, refer to Appendix F.
@@
where $K^{*}$ refers to the subset of state dimensions describing the task-relevant target, specifically the Cartesian coordinates of the end-effector position. All learnable parameters are denoted by $\bm{\theta_{\pi}}$, and we write the policy model as $\bm{\pi}(\bm{u}_{t}|\bm{s}_{t},\bm{s}_{t}^{*},\bm{\theta_{\pi}})$. See Appendix G for details.
@@
All state and action variables are linearly rescaled to [1,1][-1,1] and encoded via a learned linear projection. Specifically, the scalar input vector $\bm{x}_{t}$ ($\bm{s}_{t},\bm{u}_{t}]$ for prediction network, $\bm{s}_{t},\bm{s}^{*}_{t}]$ for policy network) is linearly transformed by
@@
where $W_{\text{in}}$ and $\bm{b}_{\text{in}}$ are learned parameters. $I_{\text{inj},t}$ serves as the input current to the first spiking layer. Injected currents are not constrained to be positive; instead, both the sign and magnitude of $I_{\text{inj},t}$ are determined by the learned input projection, allowing negative-valued inputs to be represented naturally through negative currents. Alternative encoding schemes were explored but did not yield consistent improvements and are therefore omitted from the final model.
To obtain the continuous-valued output vector $\bm{y}_{t}$ (actions or predictions), we use the membrane potentials $U_{t}$ of a non-spiking layer. These first and final weight matrices are therefore scaled differently compared to other weight parameters in the network to generate stable internal dynamics as well as output on the correct expected magnitude. The output scaling also differs for prediction and control models, as the predicted changes in state $\Delta\hat{\bm{s}}(t)$ are typically much smaller than the control signals $\bm{u}(t)$. In the policy network, a $\tanh$ activation function ensures that control outputs remain within [1,1][-1,1]. To mitigate saturation-induced gradient collapse, we apply an auxiliary $\ell_{2}$ penalty to the policy’s pre-$\tanh$ activations when they exceed the range [3,3][-3,3].